# Generating optimal Gravitational-Wave template banks with metric-preserving autoencoders

### v1

@@ 3 Conclusions
We conclude by discussing the possibility of performing dimensionality reduction using autoencoders in the space of another set of SVD coefficients, this time in the context of cosmology. More precisely, we have in mind the COBRA decomposition of the linear matter power spectrum of Ref. Bakx:2024zgu. One can generate a set of linear power spectra as a function of a number of cosmological parameters ($\omega_{\rm c}$, $\omega_{\rm b}$ and $n_{\rm s}$, for example), and construct its SVD to find a small number of basis functions with which to represent any linear power spectrum. More precisely, following Ref. Bakx:2024zgu we generate power spectra at $500$ wavenumbers between $\smash{0.8\times 10^{-4}\,{\rm Mpc}^{-1}}$ and $\smash{4\,{\rm Mpc}^{-1}}$ on a regular ($40\times 20\times 20$) grid $(\omega_{\rm c},\omega_{\rm b},n_{\rm s})\in[0.08,0.175]\times[0.020,0.025]\times[0.8,1.2]$ using CAMB (see the last column of Tab. 1 of Bakx:2024zgu for more details), and we extract the first $16$ $c^{A}$ as in Ref. Bakx:2024zgu (with the difference that besides rescaling the power spectra by the mean of the templates, we also subtract the mean of these rescaled spectra before performing the SVD): this is our training set. Then, we uniformly generate $32256$ random validation spectra in the same $\omega_{\rm c}$, $\omega_{\rm b}$ and $n_{\rm s}$, range. There are two differences with the GW case discussed in this paper. First, now we would like that the latent space is the space of cosmological parameters $(\omega_{\rm c},\omega_{\rm b},n_{\rm s})$, effectively finding the “inverse” of the Boltzmann code used to generate the templates. This is less trivial than just autoencoding the $c^{A}$ as done in this paper. Second, Ref. Bakx:2024zgu only generated templates on an uniform and structured grid, without a notion of distance between the resulting spectra or the space of the $c^{A}$, unlike e.g. what was discussed in Ref. Philcox:2020zyp in the context of cosmology. Hence, we are not interested – when comparing with Ref. Bakx:2024zgu – in constructing distance-preserving autoencoders. In Fig. 6 we show a preliminary result for what we could call the training of the decoder only: we construct a simple MLP with a $3$-dimensional input layer, $8$ hidden layers each containing $128$ neurons and a GELU activation, and a $16$-dimensional output layer, and train it to reconstruct $d^{A}=c^{A}/{\cal N}$ where $\cal N$ is defined as in Section 2.2.666Unlike Section 2.2, here $\cal N$ is defined using only training data. We refer to the caption of Fig. 6 for more details about the training. While our result can already be used as an alternative to the Radial Basis Function (RBF) interpolation of Bakx:2024zgu – we checked that it has comparable speed (e.g. we can predict our $32256$ validation spectra in $\sim 660\,{\rm ms}$ on a single core) while having slightly better precision (compare the magenta line in the lower panel of Fig. 1 of Bakx:2024zgu with the right panel of Fig. 6) – it will be interesting to check if we can construct an autoencoder for the $16$ SVD coefficients using this MLP as starting point for the decoder, and to investigate in which sense distance-preserving can be useful in cosmology. We present preliminary results in Appendix F and we leave a more detailed investigation to future work.
@@
Acknowledgments — GC acknowledges support from the Institute for Advanced Study where part of this work was carried out. We thank Lorenzo Bordin for useful comments. We thank Thomas Bakx for providing the dataset “Plin_CAMB_LCDM_Wide.dat” we used as training dataset for the COBRA autoencoder and for providing the code used in Ref. Bakx:2024zgu to get the SVD of the linear matter power spectra templates. We acknowledge use of the numpy, scipy, pytorch, zuko, scikit-learn and matplotlib packages.
@@
In the Appendix we discuss different results that were not of crucial importance for the main text, but that are nevertheless interesting to understand the geometry of our template banks. Appendix F, instead, contains some preliminary results on the COBRA autoencoder.
A Toy example: embedding a sphere
It is always instructive to consider some examples that are well understood. We start with the unit sphere $S^{2}$ isometrically embedded in $3$-dimensional Euclidean space. We generate $8192$ points on the sphere using the Fibonacci sphere algorithm 2010MatGe..42...49G, and randomly split half of them in training and validation sets. We imagine that the coordinates ${\bm{x}}=(x,y,z)$ play the role of the first $10$ SVD coefficients $c^{A}$, $A=0,\dots 9$ for GW waveforms discussed in the introduction, and that we want to describe this $2$-dimensional manifold using dimensionality reduction techniques. It is clear that, in this particular example (where by construction we do not have any hierarchy between $x$, $y$ or $z$, since we only “postulated” that these come from a SVD), a dimensionality reduction consisting in taking only $2$ of the three “SVD coefficients” (i.e. only $x$ and $y$, or only $y$ and $z$, etc.) would completely fail to capture the manifold, hence we immediately move to other methods. Inspired also here by Ref. Wadekar:2023kym and by the main text, we then train a random forest (also here we use https://scikit-learn.org/stable/modules/generated/sklearn.ensemble.RandomForestRegressor.html, with $50$ trees and maximum depth of $50$) to predict the $z$ coordinate as a function of $(x,y)$. We plot the result in the left panel of Fig. 7. We see that the random forest makes an extremely poor job of describing $S^{2}$. The reason is that for any given $(x,y,z)$ with unit norm, also (x,y,z)(x,y,{-z}) has $|{\bm{x}}|=1$, and the random forest cannot capture this. This is confirmed by the right panel of Fig. 7, where we use a random forest with the same architecture to predict $z={\rm RF}(x,y)$ but only for $z\geq 0$: the random forest manages to fit the half sphere just fine.
@@ A Toy example: embedding a sphere
and finally the Riemann tensor $\smash{R^{\rho}_{\hphantom{\rho}\sigma\mu\nu}=\partial_{\mu}\Gamma^{\rho}_{\nu\sigma}-\partial_{\nu}\Gamma^{\rho}_{\mu\sigma}+\Gamma^{\rho}_{\mu\lambda}\Gamma^{\lambda}_{\nu\sigma}-\Gamma^{\rho}_{\nu\lambda}\Gamma^{\lambda}_{\mu\sigma}}$. However, we have found it not so straightforward in pytorch to compute derivatives of a vector function – the decoder, in our case – past the second, which would be needed to compute the Riemann tensor. Hence we pursue a different route using the connection between holonomy and curvature, i.e. the non-Abelian Stokes theorem – see e.g. Appendix I of Ref. Carroll:2004st and Ref. Broda:2000id. Let us consider the parallel propagator $\smash{P^{\mu}_{\hphantom{\mu}\nu}(\lambda,\lambda_{0})}$ along a path $e^{\mu}(\lambda)$, $\lambda\in[\lambda_{0},\lambda_{1}]$. This satisfies the ODE
@@
In Fig. 9a we show the validation embeddings and a collection of $12288$ randomly-chosen points $e\in{\cal S}$. This region is well inside the domain of the decoder (notice that for this experiment we consider only the autencoder trained with $\alpha=1$, $\beta=1$ and $\gamma=5$, shown in orange in Figs. 7 and 8). Then, we compute the holonomy around circles of radius $\epsilon=10^{-3}$ centered at each of these points, [^9] We solve the ODE of Eq. (16) numerically, after refining the autoencoder to torch.float64 precision. and the integral inside the disk $D_{\epsilon}(e)$ that appears at the denominator of Eq. (18). The result is shown in Fig. 9b, where we see that the Ricci scalar is nicely distributed around the constant value $R=2$ as expected, consistently with errors coming from an inexact description of the sphere, higher orders in Eq. (17), etc.
@@
B Toy example: embedding a cylinder
@@
In this short appendix we show the results of training the same autoencoder of Appendix A with a distance preservation term in the loss different than Eq. (13). More precisely, the loss we use now is
@@
Fig. 11b is the equivalent of Fig. 8, including also the autoencoder trained with the new loss. While we see that even with this new loss the distances are not preserved as well as for the template banks discussed in Section 2.2, there is a marked improvement over the autoencoders of Appendix A, especially for small distances. It is also interesting that in this case all points seem to lie above the diagonal: while distances are not preserved, the latent-space ones seem to be consistently larger than the embedding-space ones. It would be interesting to investigate this example in more detail, possibly changing Eq. (19) to compute the Frobenius loss squared, or choosing different values for the $\alpha/\beta$ ratio.
@@ E Curvature of GW template banks
We choose $12288$ physical points for the two banks, making sure that they are at least a distance of 0.050.05131313Notice that this is intended as Euclidean distance in $2$ dimensions, consistently with the formulation of the non-Abelian Stokes theorem. from the $99.7\%$ CL contour in order to be sure that all the circles around which we compute the holonomy are also inside this contour. We then carry out the same procedure as in Appendix A to compute the numerator and denominator of Eq. (21). The results are shown in Fig. 15: we see that while both banks have a distribution of the curvature consistent with zero, the higher-mass bank has a spread in curvature which is around two orders of magnitude larger than the lower mass bank. This was expected from what we saw in Fig. 14. Also, it is important to emphasize that the curvature is an invariant of the manifold: hence this result is completely independent of the choices of $\alpha,\beta,\gamma$ we use when training the autoencoders (as long as both of them describe the respective manifolds with a comparable reconstruction error).
@@ F Autoencoder for COBRA
In Section 3 we discussed the idea of constructing an autoencoder for the SVD decomposition of the linear power spectrum of Ref. Bakx:2024zgu. In Fig. 6 we have shown what we could call the decoder in this setup. We just trained a MLP to reproduce the $c^{A}$ in terms of the cosmological parameters $\theta=(\omega_{\rm c},\omega_{\rm b},n_{\rm s})$, reaching a precision slightly better than the RBF interpolation of Bakx:2024zgu. If we want to actually obtain a full autoencoder, we also need to construct and train the encoder, with the constraint that now we want that the latent space is the space of $\theta^{\mu}$. To this purpose, we constructed an encoder with a $16$-dimensional input layer, $8$ hidden layers each containing $128$ neurons and a GELU activation, and a $3$-dimensional output layer to go from the $d^{A}$, defined as $c^{A}/{\cal N}$, to the latent space of some variables $e^{\mu}$. Then, instead of forcing via a simple $\rm MSE$} loss that $e^{\mu}\approx\theta^{\mu}$, in order to add flexibility we constructed an invertible transformation $f^{\mu}$ using zuko with the goal that $\theta^{\mu}=f^{\mu}(e)$. More precisely, $f^{\mu}$ is a composition of a sequence of $24$ Masked Autoregressive Transforms ($4$ repetitions of $3!=6$ permutations) that utilize Monotonic Affine Transforms as their univariate components. The parameters for the Monotonic Affine Transform are predicted by an internal neural network architecture consisting of $4$ hidden layers each with $64$ neurons and a GELU activation function. These transforms are followed by the composition with an untrainable sigmoid transform and finally with an affine transformation acting separately on each $e^{\mu}$ in order to bring them inside the box covered by the uniform and structured grid used in Ref. Bakx:2024zgu. In other words, our architecture is
@@
where $\smash{{\cal P}_{\theta}}$ is the uniform distribution inside the box used in Ref. Bakx:2024zgu. This could be a next step once we move on from a training set consisting of points on a uniform and structured grid (for which the distribution is more akin to a sum of Dirac delta functions). Moreover, having the function $f$ could be useful if later we want to impose the latent space to have some specific property, e.g. distance preservation, or as a starting point for the training of a variational autoencoder to learn the posterior on cosmological parameters.

### v2

@@ 5.1 Application to modeling the cosmological matter power spectrum
We then uniformly generate $32256$ random validation spectra in the same $\omega_{\rm c}$, $\omega_{\rm b}$ and $n_{\rm s}$, range. There are two differences with the GW case discussed in this paper. First, now we would like that the latent space is the space of cosmological parameters $(\omega_{\rm c},\omega_{\rm b},n_{\rm s})$, effectively finding the “inverse” of the Boltzmann code used to generate the templates. This is less trivial than just autoencoding the $c^{A}$ as done in this paper. Second, Ref. Bakx et al. (2025) only generated templates on an uniform and structured grid, without a notion of distance between the resulting spectra or the space of the $c^{A}$, unlike e.g. what was discussed in Ref. Philcox et al. (2021) in the context of cosmology. Hence, we are not interested – when comparing with Ref. Bakx et al. (2025) – in constructing metric-preserving autoencoders. In Fig. 7 we show a preliminary result for what we could call the training of the decoder only: we construct a simple MLP with a $3$-dimensional input layer, $8$ hidden layers each containing $128$ neurons and a GELU activation, and a $16$-dimensional output layer, and train it to reconstruct $d^{A}=c^{A}/{\cal N}$ where $\cal N$ is defined as in Section 3.555Unlike Section 3, here $\cal N$ is defined using only training data. We refer to the caption of Fig. 7 for more details about the training. While our result can already be used as an alternative to the Radial Basis Function (RBF) interpolation of Bakx et al. (2025) – we checked that it has comparable speed (e.g. we can predict our $32256$ validation spectra in $\sim 660\,{\rm ms}$ on a single core) while having slightly better precision (compare the magenta line in the lower panel of Fig. 1 of Bakx et al. (2025) with the right panel of Fig. 7) – it will be interesting to check if we can construct an autoencoder for the $16$ SVD coefficients using this MLP as starting point for the decoder, and to investigate in which sense a distance-preserving model can be useful in cosmology. We present preliminary results in Appendix F and we leave a more detailed investigation to future work.
6 Conclusions
Linear methods such as singular value decomposition (SVD) are used for dimensionality reduction and data compression in various domains of physics and astronomy . In this paper, we have shown that autoencoders can be used to perform more optimal dimensionality reduction than linear methods in the context of gravitational wave data analysis. Specifically, we introduce a metric-preserving autoencoder whose latent space can be used to efficiently construct gravitational wave template banks and requires us to use just two dimensions. Comparing our new template banks with a bank constructed using two SVD dimensions, our banks perform much better in the effectualness test (see Figs. 3 and 4). We also compare our results with the non-linear dimensionality-reduction method used in Ref. Wadekar et al. (2024b) (where they predict higher-order SVD coefficients from lower-order ones with random forest regressors). We find that the effectualness of our template banks is comparable to that of Wadekar et al. (2024b) across the whole mass range, while our approach allows for using a slightly smaller number of templates.
Future directions might include extending our method to include higher modes as in Wadekar et al. (2024b) or effects of spin-orbit precession. We also plan to use the autoencoder-based template banks to perform a binary black hole merger search in the data from a LIGO-Virgo-Kagra (LVK) observing run. Other than building template banks, our autoencoder method can be useful for other GW applications requiring dimensionality reduction, such as gravitational waveform modeling Field et al. (2011); Schmidt et al. (2021); Chua et al. (2019); Tiglio and Villanueva (2022); Pürrer (2014); Canizares et al. (2015); Smith et al. (2016); Field et al. (2014); Blackman et al. (2014); Herrmann et al. (2012), fast parameter estimation Messick et al. (2017), and model-independent tests of general relativity Datta et al. (2024); Saleem et al. (2022).
@@ 6 Conclusions
Acknowledgments — We thank Mark Cheung and Lorenzo Bordin for useful discussions. GC acknowledges support from the Institute for Advanced Study where part of this work was carried out. We thank Thomas Bakx for providing the dataset “Plin_CAMB_LCDM_Wide.dat” we used as training dataset for the COBRA autoencoder and for providing the code used in Ref. Bakx et al. (2025) to get the SVD of the linear matter power spectra templates. We acknowledge use of the numpy, scipy, pytorch, zuko, scikit-learn and matplotlib packages. DW is supported by NSF Grants No. AST-2307146, No. PHY-2513337, No. PHY-090003, and No. PHY-20043, by NASA Grant No. 21-ATP21-0010, by John Templeton Foundation Grant No. 62840, by the Simons Foundation [MPS-SIP-00001698, E.B.], by the Simons Foundation International [SFI-MPS-BH-00012593-02], and by Italian Ministry of Foreign Affairs and International Cooperation Grant No. PGR01167.
@@
In the Appendix we discuss different results that were not of crucial importance for the main text, but that are nevertheless interesting to understand the geometry of our template banks. Appendix F contains some preliminary results on the COBRA autoencoder used in the cosmology example in section 5.1.
A Toy example: embedding of a spherical shell
It is always instructive to consider some examples that are well understood. We start with the unit sphere $S^{2}$ isometrically embedded in $3$-dimensional Euclidean space. We generate $8192$ points on the sphere using the Fibonacci sphere algorithm González (2010), and randomly split half of them in training and validation sets. We imagine that the coordinates ${\bm{x}}=(x,y,z)$ play the role of the first $10$ SVD coefficients $c^{A}$, $A=0,\dots 9$ for GW waveforms discussed in the introduction, and that we want to describe this $2$-dimensional manifold using dimensionality reduction techniques. It is clear that, in this particular example (where by construction we do not have any hierarchy between $x$, $y$ or $z$, since we only “postulated” that these come from a SVD), a dimensionality reduction consisting in taking only $2$ of the three “SVD coefficients” (i.e. only $x$ and $y$, or only $y$ and $z$, etc.) would completely fail to capture the manifold, hence we immediately move to other methods. Inspired also here by Ref. Wadekar et al. (2024b) and by the main text, we then train a random forest (also here we use https://scikit-learn.org/stable/modules/generated/sklearn.ensemble.RandomForestRegressor.html, with $50$ trees and maximum depth of $50$) to predict the $z$ coordinate as a function of $(x,y)$. We plot the result in the left panel of Fig. 8. We see that the random forest makes an extremely poor job of describing $S^{2}$. The reason is that for any given $(x,y,z)$ with unit norm, also (x,y,z)(x,y,{-z}) has $|{\bm{x}}|=1$, and the random forest cannot capture this. This is confirmed by the right panel of Fig. 8, where we use a random forest with the same architecture to predict $z={\rm RF}(x,y)$ but only for $z\geq 0$: the random forest manages to fit the half sphere just fine.
@@ A Toy example: embedding of a spherical shell
and finally the Riemann tensor $\smash{R^{\rho}_{\hphantom{\rho}\sigma\mu\nu}=\partial_{\mu}\Gamma^{\rho}_{\nu\sigma}-\partial_{\nu}\Gamma^{\rho}_{\mu\sigma}+\Gamma^{\rho}_{\mu\lambda}\Gamma^{\lambda}_{\nu\sigma}-\Gamma^{\rho}_{\nu\lambda}\Gamma^{\lambda}_{\mu\sigma}}$. However, we have found it not so straightforward in pytorch to compute derivatives of a vector function – the decoder, in our case – past the second, which would be needed to compute the Riemann tensor. Hence we pursue a different route using the connection between holonomy and curvature, i.e. the non-Abelian Stokes theorem – see e.g. Appendix I of Ref. Carroll (2019) and Ref. Broda (2000). Let us consider the parallel propagator $\smash{P^{\mu}_{\hphantom{\mu}\nu}(\lambda,\lambda_{0})}$ along a path $e^{\mu}(\lambda)$, $\lambda\in[\lambda_{0},\lambda_{1}]$. This satisfies the ODE
@@
In Fig. 10a we show the validation embeddings and a collection of $12288$ randomly-chosen points $e\in{\cal S}$. This region is well inside the domain of the decoder (notice that for this experiment we consider only the autencoder trained with $\alpha=1$, $\beta=1$ and $\gamma=5$, shown in orange in Figs. 8 and 9). Then, we compute the holonomy around circles of radius $\epsilon=10^{-3}$ centered at each of these points, [^8] We solve the ODE of Eq. (16) numerically, after refining the autoencoder to torch.float64 precision. and the integral inside the disk $D_{\epsilon}(e)$ that appears at the denominator of Eq. (18). The result is shown in Fig. 10b, where we see that the Ricci scalar is nicely distributed around the constant value $R=2$ as expected, consistently with errors coming from an inexact description of the sphere, higher orders in Eq. (17), etc.
@@
B Toy example: embedding of a cylindrical shell
@@
In this appendix, we show the results of training the same autoencoder of Appendix A with a distance preservation term in the loss different than Eq. (13). More precisely, the loss we use now is
@@
Fig. 12b is the equivalent of Fig. 9, including also the autoencoder trained with the new loss. While we see that even with this new loss the distances are not preserved as well as for the template banks discussed in Section 3, there is a marked improvement over the autoencoders of Appendix A, especially for small distances. It is also interesting that in this case all points seem to lie above the diagonal: while distances are not preserved, the latent-space ones seem to be consistently larger than the embedding-space ones. It would be interesting to investigate this example in more detail, possibly changing Eq. (19) to compute the Frobenius loss squared, or choosing different values for the $\alpha/\beta$ ratio.
@@ E Curvature of GW template banks
We choose $12288$ physical points for the two banks, making sure that they are at least a distance of 0.050.05121212Notice that this is intended as Euclidean distance in $2$ dimensions, consistently with the formulation of the non-Abelian Stokes theorem. from the $99.7\%$ CL contour in order to be sure that all the circles around which we compute the holonomy are also inside this contour. We then carry out the same procedure as in Appendix A to compute the numerator and denominator of Eq. (21). The results are shown in Fig. 16: we see that while both banks have a distribution of the curvature consistent with zero, the higher-mass bank has a spread in curvature which is around two orders of magnitude larger than the lower mass bank. This was expected from what we saw in Fig. 15. Also, it is important to emphasize that the curvature is an invariant of the manifold: hence this result is completely independent of the choices of $\alpha,\beta,\gamma$ we use when training the autoencoders (as long as both of them describe the respective manifolds with a comparable reconstruction error).
@@ F Autoencoder for COBRA
In Section 6 we discussed the idea of constructing an autoencoder for the SVD decomposition of the linear power spectrum of Ref. Bakx et al. (2025). In Fig. 7 we have shown what we could call the decoder in this setup. We just trained a MLP to reproduce the $c^{A}$ in terms of the cosmological parameters $\theta=(\omega_{\rm c},\omega_{\rm b},n_{\rm s})$, reaching a precision slightly better than the RBF interpolation of Bakx et al. (2025). If we want to actually obtain a full autoencoder, we also need to construct and train the encoder, with the constraint that now we want that the latent space is the space of $\theta^{\mu}$. To this purpose, we constructed an encoder with a $16$-dimensional input layer, $8$ hidden layers each containing $128$ neurons and a GELU activation, and a $3$-dimensional output layer to go from the $d^{A}$, defined as $c^{A}/{\cal N}$, to the latent space of some variables $e^{\mu}$. Then, instead of forcing via a simple $\rm MSE$} loss that $e^{\mu}\approx\theta^{\mu}$, in order to add flexibility we constructed an invertible transformation $f^{\mu}$ using zuko with the goal that $\theta^{\mu}=f^{\mu}(e)$. More precisely, $f^{\mu}$ is a composition of a sequence of $24$ Masked Autoregressive Transforms ($4$ repetitions of $3!=6$ permutations) that utilize Monotonic Affine Transforms as their univariate components. The parameters for the Monotonic Affine Transform are predicted by an internal neural network architecture consisting of $4$ hidden layers each with $64$ neurons and a GELU activation function. These transforms are followed by the composition with an untrainable sigmoid transform and finally with an affine transformation acting separately on each $e^{\mu}$ in order to bring them inside the box covered by the uniform and structured grid used in Ref. Bakx et al. (2025). In other words, our architecture is
@@
where $\smash{{\cal P}_{\theta}}$ is the uniform distribution inside the box used in Ref. Bakx et al. (2025). This could be a next step once we move on from a training set consisting of points on a uniform and structured grid (for which the distribution is more akin to a sum of Dirac delta functions). Moreover, having the function $f$ could be useful if later we want to impose the latent space to have some specific property, e.g. distance preservation, or as a starting point for the training of a variational autoencoder to learn the posterior on cosmological parameters.